# On lower bounds for the distances between APN functions

### v1

@@ Abstract
Whether two distinct APN functions can have a Hamming distance of $1$ remains an open problem. In 2020, L. Budaghyan et al. introduced a new CCZ-invariant $\Pi_{F}$ which can be used to provide lower bounds on the Hamming distance between a given APN function $F\colon\mathbb{F}_{2}^{n}\to\mathbb{F}_{2}^{n}$ and other APN functions. Lower bounds on the distance from an APN function $F$ to any other are known for almost bent (AB) functions and when $F$ is a 3-to-1 quadratic function with $n$ even. In this paper, we reinterpret $\Pi_{F}$ in terms of the exclude multiplicities of the graph $\mathcal{G}_{F}=\left\{(x,F(x)):x\in\mathbb{F}_{2}^{n}\right\}$ of $F$ as a Sidon set. We establish lower bounds on the distance for even $n$ when $F$ is plateaued APN, generalize the known lower bounds for quadratic 33-to-11 function to all 3-to-1 plateaued functions (e.g. Kasami functions), and derive new lower bounds for when $F$ is the APN inverse function over $\mathbb{F}_{2^{n}}$ for $n$ odd. We also study how the exclude multiplicities of $\mathcal{G}_{F}$ are directly connected to the existence of linear structures of $\gamma_{F}$ when $F$ is plateaued and APN and the ortho-derivative when $F$ is a quadratic APN function. We also use the CCZ-invariance of exclude multiplicities to prove that the Brinkmann-Leander-Edel-Pott function is not CCZ-equivalent to a plateaued function.
@@ 1 Introduction
where $e_{\min}(\mathcal{G}_{F})=\min_{(a,b)\in(\mathbb{F}_{2}^{n})^{2}\setminus\mathcal{G}_{F}}\operatorname{mult}_{\mathcal{G}_{F}}(a,b)$. The proofs of each of our main results all rely on finding a lower bound on $e_{\min}(\mathcal{G}_{F})$, and in some cases, we determine the exact exclude multiplicities of $\mathcal{G}_{F}$.
@@
In Section 4, we study the exclude multiplicities of the graphs of plateaued APN functions, and we prove Theorems 1.1 and 1.2. Since all plateaued APN functions are AB when $n$ is odd, we mostly focus on the case when $n$ is even. We express the exclude multiplicities of $\mathcal{G}_{F}$ in terms of a sum involving the amplitudes of the components of $F$. Moreover, we generalize a result of [5] to all plateaued functions, showing that $\Pi_{F}$ is completely determined by $\Pi_{F}^{0}$ (see Section 3) when $F$ is plateaued and APN. As a corollary, we prove that the Brinkmann-Leander-Edel-Pott function (the only known instance of an APN function not CCZ-equivalent to a power function or a quadratic function) is also not CCZ-equivalent to a plateaued function. In Section 4.4, we study quadratic APN functions with a focus on how their ortho-derivatives can be used to express exclude multiplicities. By using a recently proved fact about the algebraic degree of an ortho-derivative, we show that all exclude multiplicities of $\mathcal{G}_{F}$ are odd when $F$ is quadratic APN. As an equivalent result, we prove that ${\mathcal{B}({F})}=\left\{v\in\mathbb{F}_{2}^{n}:v\cdot F\text{ is bent}\right\}$ always has an intersection of even size with any affine hyperplane of $\mathbb{F}_{2}^{n}$ when $F$ is quadratic APN.
@@
In Section 7, we study how exclude multiplicity is related to the $\gamma_{F}$ Boolean function associated to an APN function $F$. In particular, we show that if $F$ is plateaued APN, then $\gamma_{F}$ has no nontrivial linear structures if and only if $\mathcal{G}_{F}$ does not have a point of exclude multiplicity $\frac{2^{n}-1}{3}$ (the largest possible). As a consequence, we prove that $\gamma_{F}$ has no nontrivial linear structures when $F$ is plateaued 3-to-1. We also show that $\gamma_{F}$ has no nontrivial linear structures when $F$ is quadratic APN with an ortho-derivative of nonzero nonlinearity. To conclude the paper, we list open problems in Section 8.
@@ 2 Preliminaries
For a vectorial Boolean function $F\colon\mathbb{F}_{2}^{n}\to\mathbb{F}_{2}^{n}$, the Walsh transform of $F$ is defined as $W_{F}(u,v)=\sum_{x\in\mathbb{F}_{2}^{n}}(-1)^{x\cdot u+F(x)+v}$, where $\cdot$ is the standard inner product on $\mathbb{F}_{2}^{n}$. It is a classical result of [16] that $F$ is APN if and only if $\sum_{(u,v)\in(\mathbb{F}_{2}^{n})^{2}}W_{F}^{4}(u,v)=2^{2n}(3\cdot 2^{2n}-2^{n+1})$. We say that the linearity of $F$ is the value $\mathcal{L}(F)=\max_{(u,v)\in(\mathbb{F}_{2}^{n})^{2}\setminus(0,0)}|W_{F}(u,v)|$, and the nonlinearity of $F$ is $\operatorname{\mathcal{NL}}(F)=2^{n-1}-\frac{1}{2}\mathcal{L}(F)$. Oftentimes, we identify $\mathbb{F}_{2}^{n}$ with $\mathbb{F}_{2^{n}}$, and in which case, we define $u\cdot v$ as $\operatorname{Tr}(uv)$ where $\operatorname{Tr}$ is the absolute trace function $\operatorname{Tr}(x)=\sum_{i=0}^{n-1}x^{2^{i}}$ from $\mathbb{F}_{2^{n}}$ onto $\mathbb{F}_{2}$. The Walsh transform of $F$ can also be expressed as the Fourier-Hadamard transform of the indicator $1_{\mathcal{G}_{F}}$ of $\mathcal{G}_{F}$. Recall that the Fourier-Hadamard transform of a function $\varphi\colon\mathbb{F}_{2}^{n}\to\mathbb{Z}$ is the function $\widehat{\varphi}(u)=\sum_{x\in\mathbb{F}_{2}^{n}}(-1)^{x\cdot u}\varphi(x)$. We will also use the convolutional product, which for any two functions $\varphi,\psi\colon\mathbb{F}_{2}^{n}\to\mathbb{Z}$ is given by $(\varphi\times\psi)(u)=\sum_{x\in\mathbb{F}_{2}^{n}}\varphi(x)\psi(x+u)$.
@@
APN functions also have a natural connection to Sidon sets as the graph of $F$ is a Sidon set in $(\mathbb{F}_{2}^{n})^{2}$ if and only if $F$ is APN. A Sidon set is a set $S\subseteq\mathbb{F}_{2}^{n}$ such that $x+y+z+w=0$ has no solutions for all pairwise distinct $x,y,z,w\in S$. A Sidon set is called maximal if it is not contained in any strictly larger Sidon set. In [12] used the connection of APN functions and Sidon sets to prove that 2.1 is equivalent to the following conjecture.
@@ 3 Distance between APN functions
We now compute the size of $\Pi_{F}^{\beta}(b)$ in terms of $\gamma_{F}$, and this will be useful in demonstrating the connection between shifted derivatives and exclude multiplicity. Note that for any $u\in\mathbb{F}_{2}^{n}$, we write the Kronecker delta as $\delta_{0}(u)$, which takes value $1$ if and only if $u=0$ and value 0 otherwise.
@@
Since $\widehat{\varphi\otimes\psi}=\widehat{\varphi}\widehat{\psi}$ for any functions $\varphi,\psi\colon\mathbb{F}_{2}^{n}\to\mathbb{Z}$ (see [11, Proposition 11]) and using the fact that $\widehat{1_{\mathcal{G}_{F}}}=W_{F}$, the Fourier-Hadamard transform of $(1_{\mathcal{G}_{F}}\otimes 1_{\mathcal{G}_{F}})\otimes 1_{\mathcal{G}_{F}}$ is $\widehat{(1_{\mathcal{G}_{F}}\otimes 1_{\mathcal{G}_{F}})}\widehat{1_{\mathcal{G}_{F}}}=(\widehat{1_{\mathcal{G}_{F}}}\widehat{1_{\mathcal{G}_{F}}})\widehat{1_{\mathcal{G}_{F}}}=W_{F}^{3}$. Therefore, taking the Fourier-Hadamard transform of both sides, we have $(1_{\mathcal{G}_{F}}\otimes 1_{\mathcal{G}_{F}})\otimes 1_{\mathcal{G}_{F}}=\frac{1}{2^{2n}}\widehat{W_{F}^{3}}$, and so
@@
In particular, when $(\beta,b)\notin\mathcal{G}_{F}$ and $F$ is APN, we have
@@
For any APN function $F$, any exclude point of $\mathcal{G}_{F}$ has exclude multiplicity at most $\lfloor\frac{2^{n}}{3}\rfloor$ (see [19, Corollary 3.4]). Therefore, $m_{F}=\min\left\{3e_{\min}(\mathcal{G}_{F}),2^{n}\right\}=3e_{\min}(\mathcal{G}_{F})$, and so the bound of Corollary 3.5 is equivalent to the bound of Lemma 3.1.
@@ 4.2 Properties of the exclude multiplicities of $\mathcal{G}_{F}$ when $F$ is plateaued APN
Theorem 4.3 is very useful for classifying plateaued functions, and in particular, we know that the size of the set
@@
is invariant of the choice of $x\in\mathbb{F}_{2}^{n}$ when $w\in\mathbb{F}_{2}^{n}$ is fixed and $F\colon\mathbb{F}_{2}^{n}\to\mathbb{F}_{2}^{n}$ is a plateaued function. By replacing $b$ with $x+b$, the size of the above set is the same as
@@
In the particular case that $F$ is also APN, then $\delta_{F}=2\gamma_{F}+2^{n}\delta_{(0,0)}$. So when $F$ is APN and $w\neq 0$, we have $\sum_{a\in\mathbb{F}_{2}^{n}}\delta_{F}(a+x,F(a)+F(x)+w)=2\sum_{a\in\mathbb{F}_{2}^{n}}\gamma_{F}(a+x,F(a)+F(x)+w)$, and by Proposition 3.3 we know that $\sum_{a\in\mathbb{F}_{2}^{n}}\gamma_{F}(a+x,F(a)+F(x)+w)=3\operatorname{mult}_{\mathcal{G}_{F}}(x,F(x)+w)$. Hence
@@
when $F$ is APN and $w\neq 0$. So Theorem 4.3 implies that if $F$ is plateaued APN and $w\neq 0$, then $\operatorname{mult}_{\mathcal{G}_{F}}(x,F(x)+w)$ does not depend on $x\in\mathbb{F}_{2}^{n}$. For $F\colon\mathbb{F}_{2}^{n}\to\mathbb{F}_{2}^{n}$ quadratic, it was shown in [5, Proposition 5] that $\Pi_{F}^{\beta}$ does not depend on $\beta\in\mathbb{F}_{2}^{n}$, and so if $F$ is APN quadratic, the exclude multiplicities $\operatorname{mult}_{\mathcal{G}_{F}}(a,b)$ as $b$ ranges across $\mathbb{F}_{2}^{n}\setminus\left\{a\right\}$ does not depend on the choice of $a\in\mathbb{F}_{2}^{n}$ by Proposition 3.3. In the following proposition, we prove this result for all plateaued APN functions.
@@
For any non-negative integer $k$, let $z_{k}=|\left\{b\in\mathbb{F}_{2}^{n}\setminus\left\{F(0)\right\}:\operatorname{mult}_{\mathcal{G}_{F}}(0,b)=k\right\}|$. By Proposition 4.4, we know that for any $a\in\mathbb{F}_{2}^{n}$, there exists a bijection from $\left\{a\right\}\times(\mathbb{F}_{2}^{n}\setminus\left\{F(a)\right\})$ to $\left\{0\right\}\times(\mathbb{F}_{2}^{n}\setminus\left\{F(0)\right\})$ that preserves exclude multiplicity. Therefore, $m_{k}=2^{n}z_{k}$ for all $k\geq 0$, so $2^{n}|m_{k}$. ∎
From this, we prove that the Brinkmann-Leander-Edel-Pott function $F\colon\mathbb{F}_{2^{6}}\to\mathbb{F}_{2^{6}}$ (see [23], [11, Section 11.5.3], or [4]), given by
@@
where $u\in\mathbb{F}_{2^{n}}$ is primitive, cannot be CCZ-equivalent to a plateaued function. This improves our understanding on this sporadic APN function as the Brinkmann-Leander-Edel-Pott function was only previously known to not be CCZ-equivalent to a quadratic function or monomial function [23].
@@
Our proof technique for showing that the Brinkmann-Leander-Edel-Pott function is not CCZ-equivalent to a plateaued function also yields a practical test that can be applied to newly discovered APN functions in the future.
@@
By using the following proposition, we will be able to rephrase minimizing exclude multiplicity in terms of minimizing the sum of squared amplitudes in any linear hyperplane as the sum $\sum_{v\in\mathbb{F}_{2}^{n}}\lambda_{v}^{2}$ is equal to $2^{n}(3\cdot 2^{n}-2)$ precisely when $F$ is APN.
Proposition 4.8 ([10, Proposition 9]).
@@
In particular, if $F$ is APN, then eq. 5 is an equality.
@@
Let $F\colon\mathbb{F}_{2}^{n}\to\mathbb{F}_{2}^{n}$ be a plateaued APN function, and let $\lambda_{v}$ denote the amplitude of $v\cdot F$ for all $v\in\mathbb{F}_{2}^{n}$. If for a non-negative integer $k$, we have
$\sum_{v\in\left\{c\right\}^{\perp}}\lambda_{v}^{2}\geq 3\cdot 2^{n}k+2^{n-1}(3\cdot 2^{n}-2)$ (6)
for all nonzero $c\in\mathbb{F}_{2}^{n}$, then $e_{\min}(\mathcal{G}_{F})\geq k$.
@@
Suppose $k$ is a non-negative integer such that for all nonzero $c\in\mathbb{F}_{2}^{n}$ eq. 6 holds. For any $(a,b)\in(\mathbb{F}_{2}^{n})^{2}\setminus\mathcal{G}_{F}$, we have by Corollary 4.9 that
$\operatorname{mult}_{\mathcal{G}_{F}}(a,b)$ $\geq\frac{1}{3\cdot 2^{n}}\left(3\cdot 2^{n}k+2^{n-1}(3\cdot 2^{n}-2)-2^{n-1}(3\cdot 2^{n}-2)\right)=k.$
So $e_{\min}(\mathcal{G}_{F})\geq k$. ∎
@@
In [34, Proof of Theorem 3], it was shown that there cannot exist a $(\frac{n}{2}+1)$-dimensional linear subspace $T$ of $\mathbb{F}_{2}^{n}$ such that $T\setminus\left\{0\right\}$ is contained in $\mathcal{B}({F})$} (otherwise we could easily construct a function $\mathbb{F}_{2}^{n}\to\mathbb{F}_{2}^{n/2+1}$ whose component functions are all bent, but this is impossible by Nyberg’s bound). Hence, $S\cap{\mathcal{B}({F})}$ cannot contain all nonzero elements of a linear subspace of dimension $\frac{n}{2}+1$ because it is a subset of $\mathcal{B}({F})$}. So $S\setminus({\mathcal{B}({F})}\cup\left\{0\right\})$ must intersect every linear subspace of $S$ of dimension $\frac{n}{2}+1$. Also, note that $\frac{n}{2}+1=d+1-(d-\frac{n}{2})$. So, by applying a linear isomorphism from $S$ to $\mathbb{F}_{2}^{d}$ and considering the image of $S\setminus({\mathcal{B}({F})}\cup\left\{0\right\})$, it follows from Corollary 4.11 that
@@
From this, we are able to derive lower bounds on the exclude multiplicities of the graphs of APN functions (for $n$ even).
@@
Assume $n\geq 4$ is even. Let $F\colon\mathbb{F}_{2}^{n}\to\mathbb{F}_{2}^{n}$ be a plateaued APN function. Then $\operatorname{mult}_{\mathcal{G}_{F}}(a,b)\geq 2^{\frac{n}{2}-1}-1$ for all $(a,b)\in(\mathbb{F}_{2}^{n})^{2}\setminus\mathcal{G}_{F}$.
@@
Let $c=F(a)+b$ for $(a,b)\in(\mathbb{F}_{2}^{n})^{2}\setminus\mathcal{G}_{F}$. Let $N=|\left\{c\right\}^{\perp}\setminus({\mathcal{B}({F})}\cup\left\{0\right\})|$. By applying Corollary 4.9, we have

### v2

@@ Abstract
Whether two distinct APN functions can have a Hamming distance of $1$ remains an open problem. In 2020, L. Budaghyan et al. introduced a new CCZ-invariant $\Pi_{F}$ which can be used to provide lower bounds on the Hamming distance between a given APN function $F\colon\mathbb{F}_{2}^{n}\to\mathbb{F}_{2}^{n}$ and other APN functions. Lower bounds on the distance from an APN function $F$ to any other are known for almost bent (AB) functions and when $F$ is a 3-to-1 quadratic function with $n$ even. In this paper, we reinterpret $\Pi_{F}$ in terms of the exclude multiplicities of the graph $\mathcal{G}_{F}=\{(x,F(x)):x\in\mathbb{F}_{2}^{n}\}$ of $F$ as a Sidon set. We establish lower bounds on the distance for even $n$ when $F$ is plateaued APN, generalize the known lower bounds for quadratic 33-to-11 function to all 3-to-1 plateaued functions (e.g. Kasami functions), and derive new lower bounds for when $F$ is the APN inverse function over $\mathbb{F}_{2^{n}}$ for $n$ odd. We also study how the exclude multiplicities of $\mathcal{G}_{F}$ are directly connected to the existence of linear structures of $\gamma_{F}$ when $F$ is plateaued and APN and the ortho-derivative when $F$ is a quadratic APN function. We also use the CCZ-invariance of exclude multiplicities to prove that the Brinkmann-Leander-Edel-Pott function is not CCZ-equivalent to a plateaued function.
@@ 1 Introduction
where $e_{\min}(\mathcal{G}_{F})=\min_{(a,b)\in(\mathbb{F}_{2}^{n})^{2}\setminus\mathcal{G}_{F}}\operatorname{mult}_{\mathcal{G}_{F}}(a,b)$. The proofs of each of our main results all rely on finding a lower bound on $e_{\min}(\mathcal{G}_{F})$, and in some cases, we determine the exact exclude multiplicities of $\mathcal{G}_{F}$. The lower bounds on the distances for different families of APN functions are summarized in Table 1.
@@
In Section 4, we study the exclude multiplicities of the graphs of plateaued APN functions, and we prove Theorems 1.1 and 1.2. Since all plateaued APN functions are AB when $n$ is odd, we mostly focus on the case when $n$ is even. We express the exclude multiplicities of $\mathcal{G}_{F}$ in terms of a sum involving the amplitudes of the components of $F$. Moreover, we generalize a result of [6] to all plateaued functions, showing that $\Pi_{F}$ is completely determined by $\Pi_{F}^{0}$ (see Section 3) when $F$ is plateaued and APN. As a corollary, we prove that the Brinkmann-Leander-Edel-Pott function (the only known instance of an APN function not CCZ-equivalent to a power function or a quadratic function) is also not CCZ-equivalent to a plateaued function. We also prove that the exclude multiplicities of $\mathcal{G}_{F}$ are odd when $F$ is plateaued APN ($n\geq 3$) via a characterization in terms of the intersections of ${\mathcal{B}({F})}=\left\{v\in\mathbb{F}_{2}^{n}:v\cdot F\text{ is bent}\right\}$ with affine hyperplanes. In Section 4.4, we study quadratic APN functions with a focus on how their ortho-derivatives can be used to express exclude multiplicities.
@@
In Section 7, we study how exclude multiplicity is related to the $\gamma_{F}$ Boolean function associated to an APN function $F$. In particular, we show that if $F$ is plateaued APN, then $\gamma_{F}$ has no nontrivial linear structures if and only if $\mathcal{G}_{F}$ does not have a point of exclude multiplicity $\frac{2^{n}-1}{3}$ (the largest possible), and we show that this latter condition always holds, proving that $\gamma_{F}$ has no nontrivial linear structures when $F$ is plateaued APN. To conclude the paper, we list open problems in Section 8.
@@ 2 Preliminaries
For a vectorial Boolean function $F\colon\mathbb{F}_{2}^{n}\to\mathbb{F}_{2}^{n}$, the Walsh transform of $F$ is defined as $W_{F}(u,v)=\sum_{x\in\mathbb{F}_{2}^{n}}(-1)^{x\cdot u+F(x)+v}$, where $\cdot$ is the standard inner product on $\mathbb{F}_{2}^{n}$. It is a classical result of [17] that $F$ is APN if and only if $\sum_{(u,v)\in(\mathbb{F}_{2}^{n})^{2}}W_{F}^{4}(u,v)=2^{2n}(3\cdot 2^{2n}-2^{n+1})$. We say that the linearity of $F$ is the value $\mathcal{L}(F)=\max_{(u,v)\in(\mathbb{F}_{2}^{n})^{2}\setminus(0,0)}|W_{F}(u,v)|$, and the nonlinearity of $F$ is $\operatorname{\mathcal{NL}}(F)=2^{n-1}-\frac{1}{2}\mathcal{L}(F)$. Oftentimes, we identify $\mathbb{F}_{2}^{n}$ with $\mathbb{F}_{2^{n}}$, and in which case, we define $u\cdot v$ as $\operatorname{Tr}(uv)$ where $\operatorname{Tr}$ is the absolute trace function $\operatorname{Tr}(x)=\sum_{i=0}^{n-1}x^{2^{i}}$ from $\mathbb{F}_{2^{n}}$ onto $\mathbb{F}_{2}$. The Walsh transform of $F$ can also be expressed as the Fourier-Hadamard transform of the indicator $1_{\mathcal{G}_{F}}$ of $\mathcal{G}_{F}$. Recall that the Fourier-Hadamard transform of a function $\varphi\colon\mathbb{F}_{2}^{n}\to\mathbb{Z}$ is the function $\widehat{\varphi}(u)=\sum_{x\in\mathbb{F}_{2}^{n}}(-1)^{x\cdot u}\varphi(x)$. We will also use the convolutional product, which for any two functions $\varphi,\psi\colon\mathbb{F}_{2}^{n}\to\mathbb{Z}$ is given by $(\varphi\otimes\psi)(u)=\sum_{x\in\mathbb{F}_{2}^{n}}\varphi(x)\psi(x+u)$.
@@
APN functions also have a natural connection to Sidon sets as the graph of $F$ is a Sidon set in $(\mathbb{F}_{2}^{n})^{2}$ if and only if $F$ is APN. A Sidon set is a set $S\subseteq\mathbb{F}_{2}^{n}$ such that $x+y+z+w=0$ has no solutions for all pairwise distinct $x,y,z,w\in S$. A Sidon set is called maximal if it is not contained in any strictly larger Sidon set. In [16], Carlet used the connection of APN functions and Sidon sets to prove that 2.1 is equivalent to the following conjecture.
@@ 3 Distance between APN functions
We now compute the size of $\Pi_{F}^{\beta}(b)$ in terms of $\gamma_{F}$, and this will be useful in demonstrating the connection between shifted derivatives and exclude multiplicity. Note that for any $u\in\mathbb{F}_{2}^{n}$, we write the Dirac delta as $\delta_{0}(u)$, which takes value $1$ if and only if $u=0$ and value 0 otherwise.
@@
Since $\widehat{\varphi\otimes\psi}=\widehat{\varphi}\widehat{\psi}$ for any functions $\varphi,\psi\colon\mathbb{F}_{2}^{n}\to\mathbb{Z}$ (see [14, Proposition 11]) and using the fact that $\widehat{1_{\mathcal{G}_{F}}}=W_{F}$, the Fourier-Hadamard transform of $(1_{\mathcal{G}_{F}}\otimes 1_{\mathcal{G}_{F}})\otimes 1_{\mathcal{G}_{F}}$ is $\widehat{(1_{\mathcal{G}_{F}}\otimes 1_{\mathcal{G}_{F}})}\widehat{1_{\mathcal{G}_{F}}}=(\widehat{1_{\mathcal{G}_{F}}}\widehat{1_{\mathcal{G}_{F}}})\widehat{1_{\mathcal{G}_{F}}}=W_{F}^{3}$. Therefore, taking the Fourier-Hadamard transform of both sides, we have $(1_{\mathcal{G}_{F}}\otimes 1_{\mathcal{G}_{F}})\otimes 1_{\mathcal{G}_{F}}=\frac{1}{2^{2n}}\widehat{W_{F}^{3}}$, and so
@@
In particular, when $(\beta,b)\notin\mathcal{G}_{F}$, we have
@@
For any APN function $F$, any exclude point of $\mathcal{G}_{F}$ has exclude multiplicity at most $\lfloor\frac{2^{n}}{3}\rfloor$ (see [20, Corollary 3.4]). Therefore, $m_{F}=\min\left\{3e_{\min}(\mathcal{G}_{F}),2^{n}\right\}=3e_{\min}(\mathcal{G}_{F})$, and so the bound of Corollary 3.5 is equivalent to the bound of Lemma 3.1.
@@ 4.2 Properties of the exclude multiplicities of $\mathcal{G}_{F}$ when $F$ is plateaued APN
Hence, for a plateaued function $F\colon\mathbb{F}_{2}^{n}\to\mathbb{F}_{2}^{n}$, the size of the set
@@
is invariant of the choice of $x\in\mathbb{F}_{2}^{n}$ when $w\in\mathbb{F}_{2}^{n}$ is fixed. By replacing $b$ with $x+b$, the size of the above set is the same as
@@
Let us consider the case that $F$ is APN. Then $\delta_{F}=2\gamma_{F}+2^{n}\delta_{(0,0)}$. So, if $w\neq 0$, then $\sum_{a\in\mathbb{F}_{2}^{n}}\delta_{F}(a+x,F(a)+F(x)+w)=2\sum_{a\in\mathbb{F}_{2}^{n}}\gamma_{F}(a+x,F(a)+F(x)+w)$, and by Proposition 3.3 we know that $\sum_{a\in\mathbb{F}_{2}^{n}}\gamma_{F}(a+x,F(a)+F(x)+w)=3\operatorname{mult}_{\mathcal{G}_{F}}(x,F(x)+w)$. Hence
@@
for any $w\neq 0$. So Theorem 4.3 implies that if $F$ is plateaued and $w\neq 0$, then $\operatorname{mult}_{\mathcal{G}_{F}}(x,F(x)+w)$ does not depend on $x\in\mathbb{F}_{2}^{n}$. For $F\colon\mathbb{F}_{2}^{n}\to\mathbb{F}_{2}^{n}$ quadratic, it was shown in [6, Proposition 5] that $\Pi_{F}^{\beta}$ does not depend on $\beta\in\mathbb{F}_{2}^{n}$, and as a consequence, the exclude multiplicities $\operatorname{mult}_{\mathcal{G}_{F}}(a,b)$ as $b$ ranges across $\mathbb{F}_{2}^{n}\setminus\left\{F(a)\right\}$ does not depend on the choice of $a\in\mathbb{F}_{2}^{n}$ by Proposition 3.3. In the following proposition, we prove this result for all plateaued APN functions.
@@
For any non-negative integer $k$, let $z_{k}=|\left\{b\in\mathbb{F}_{2}^{n}\setminus\left\{F(0)\right\}:\operatorname{mult}_{\mathcal{G}_{F}}(0,b)=k\right\}|$. By Proposition 4.4, we know that for any $a\in\mathbb{F}_{2}^{n}$ there exists a bijection from $\left\{a\right\}\times(\mathbb{F}_{2}^{n}\setminus\left\{F(a)\right\})$ to $\left\{0\right\}\times(\mathbb{F}_{2}^{n}\setminus\left\{F(0)\right\})$ that preserves exclude multiplicity. Therefore, $m_{k}=2^{n}z_{k}$ for all $k\geq 0$, so $2^{n}|m_{k}$. ∎
From this, we prove that the Brinkmann-Leander-Edel-Pott function $F\colon\mathbb{F}_{2^{6}}\to\mathbb{F}_{2^{6}}$ (see [24], [14, Section 11.5.3], or [5]), given by
@@
where $u\in\mathbb{F}_{2^{n}}$ is primitive, cannot be CCZ-equivalent to a plateaued function. This improves our understanding of this sporadic APN function as the Brinkmann-Leander-Edel-Pott function was only previously known to not be CCZ-equivalent to a quadratic function or monomial function [24].
@@
Our proof technique for showing that the Brinkmann-Leander-Edel-Pott function is not CCZ-equivalent to a plateaued function also yields a practical test that can be applied to newly discovered APN functions.
@@
By using the following proposition, we are able to rephrase minimizing exclude multiplicity in terms of minimizing the sum of squared amplitudes in any linear hyperplane. This is because $\sum_{v\in\mathbb{F}_{2}^{n}}\lambda_{v}^{2}=2^{n}(3\cdot 2^{n}-2)$ precisely when $F$ is APN.
Proposition 4.8 ([13, Proposition 9]).
@@
In particular, if $F$ is APN, then (5) is an equality.
@@
Let $F\colon\mathbb{F}_{2}^{n}\to\mathbb{F}_{2}^{n}$ be a plateaued APN function, and let $\lambda_{v}$ denote the amplitude of $v\cdot F$ for all $v\in\mathbb{F}_{2}^{n}$. For any $(a,b)\in(\mathbb{F}_{2}^{n})^{2}\setminus\mathcal{G}_{F}$, we have
$\sum_{v\in\left\{b+F(a)\right\}^{\perp}}\lambda_{v}^{2}=3\cdot 2^{n}\operatorname{mult}_{\mathcal{G}_{F}}(a,b)+2^{n-1}(3\cdot 2^{n}-2).$ (6)
@@
Immediate upon rearrangement of the equation from Corollary 4.9. ∎
@@
In [35, Proof of Theorem 3], it was shown that there cannot exist a $(\frac{n}{2}+1)$-dimensional linear subspace $T$ of $\mathbb{F}_{2}^{n}$ such that $T\setminus\left\{0\right\}$ is contained in $\mathcal{B}({F})$} (otherwise we could easily construct a function $\mathbb{F}_{2}^{n}\to\mathbb{F}_{2}^{n/2+1}$ whose component functions are all bent, but this is impossible by Nyberg’s bound). Hence, $S\cap{\mathcal{B}({F})}$ cannot contain all nonzero elements of a linear subspace of dimension $\frac{n}{2}+1$ because it is a subset of $\mathcal{B}({F})$}. So $S\setminus({\mathcal{B}({F})}\cup\left\{0\right\})$ must intersect every linear subspace of $S$ of dimension $\frac{n}{2}+1$. Also, note that $\frac{n}{2}+1=d+1-(d-\frac{n}{2})$. So, by applying a linear isomorphism from $S$ to $\mathbb{F}_{2}^{d}$ and considering the image of $S\setminus({\mathcal{B}({F})}\cup\left\{0\right\})$, it follows from Corollary 4.11 that
@@
From this, we are able to derive lower bounds on the exclude multiplicities of the graphs of plateaued APN functions (for $n$ even).
@@
Assume $n$ is even. Let $F\colon\mathbb{F}_{2}^{n}\to\mathbb{F}_{2}^{n}$ be a plateaued APN function. Then $\operatorname{mult}_{\mathcal{G}_{F}}(a,b)\geq 2^{\frac{n}{2}-1}-1$ for all $(a,b)\in(\mathbb{F}_{2}^{n})^{2}\setminus\mathcal{G}_{F}$.
@@
Let $c=F(a)+b$ for $(a,b)\in(\mathbb{F}_{2}^{n})^{2}\setminus\mathcal{G}_{F}$. Let $N=|\left\{c\right\}^{\perp}\setminus({\mathcal{B}({F})}\cup\left\{0\right\})|$. Recall that if $v\in{\mathcal{B}({F})}$, then $\lambda_{v}=2^{\frac{n}{2}}$, and otherwise, $\lambda_{v}\geq 2^{\frac{n}{2}+1}$. So